# Spectuner: A Framework for Automated Line Identification of Interstellar Molecules

### v1

@@ 4 Applying to line surveys
In this study, for speed reason, we downsample all observed spectra by averaging every two adjacent channels. The downsampling is repeated three times. Accordingly, the number of channels in the observed spectra is reduced by a factor of eight, and the RMS noise values are divided by $\sqrt{8}$ before being input into the peak finding algorithm. Secondly, for simplicity, this work neglects any absorption in the observed spectra. Given that the baselines of all the observed spectra used in this work have been calibrated to zero by the original studies, our correction is straightforward. Specifically, any channels displaying a negative value are reset to zero. In addition, the continuum spectra are not taken into account in the fitting processes.
In addition, our algorithm requires a molecule list as input. Since this work focuses on whether our method can reproduce the identification results of the previous studies rather than identifying new molecules, we use the molecules identified in previous studies as input. For practical use, our code provides a list of commonly observed molecules, e.g. methanol (CH3OH), methyl formate (CH3OCHO) and cyanoacetylene (HCCCN). Users can customize the molecule list to suit their specific applications.
@@ 4.1 Metrics
In this work, we use recall to compare our identification results with the published line survey data. The recall of this work is computed as follows. The denominator is the number of relevant lines, i.e. lines both in the line table of the test data and identified by our peak finding algorithm. When counting the relevant lines, the following points are taken into account:
@@
    If a frequency is labeled by multiple species, the corresponding lines are excluded.
@@
    A peak can be comprised of multiple spectral lines from the same molecule. In this case, the lines are merged into one.
@@
    The line surveys may use different spectral line databases, and to be consistent, we only include lines that exist in the spectral line database used in this work.
@@
    Radio recombination lines are excluded.
The numerator of the recall is the number of lines labeled by our identification algorithm that matches the test data.
@@ 4.2 Results
Table 2 shows the overall recall obtained from the test data. The overall recalls achieved using the peak matching loss are 84% - 98%, consistently higher than using the $\chi^{2}$ loss.
We illustrate the issue with using the $\chi^{2}$ loss function through an example. In Figure 4, we present the fitting results of C2H3CN during the individual fitting phase for SrgB2N-IRAM. The results based on the peak matching and $\chi^{2}$ loss functions are represented by red and blue lines respectively. It is evident that the red line gives a better result than the blue line. However, upon calculating the $\chi^{2}$ and peak matching loss values for both fits, we find that the $\chi^{2}$ value of the blue line ($\chi^{2}$ = 12780.4) is smaller than the red line ($\chi^{2}$ = 12780.9). In other words, if the $\chi^{2}$ loss is adopted, the algorithm will consider the blue line as a better result than the red line, irrespective of optimization methods. Such result is unexpected. In contrast, the proposed peak matching loss assigns a lower value to the red line ($\mathcal{L}_{\text{PM}}=-2.79$) than the blue line ($\mathcal{L}_{\text{PM}}=1.40$), which appropriately reflects the variance between the observed and fitting spectra.
@@ 4.2.1 W51-Mopra
We illustrate the best-fitting spectrum for W51-Mopra in Figure 5, and show the recalls in Figure 6. When adopting the peak matching loss function, our algorithm attains an overall recall of 91%, 13% better than using the $\chi_{2}$ loss function. The recall of methanol stands relatively low, at 62%. Our analysis suggests that the spectrum of methanol can be better described by more than one component. As demonstrated in Figure 7, by adding an additional component, the best-fitting spectrum of methanol is improved, and the recall increases to 96%.
@@ 4.2.2 OrionKL-GBT
For OrionKL-GBT, the best-fitting spectrum and the recalls of the molecules are demonstrated in Figure 8 and 9 respectively. Frayer et al. (2015) only identified lines that are brighter than 1 K, which is significantly higher than the RMS noise. Therefore, we obtain the best results on OrionKL-GBT, achieving a recall of 98 %. Our algorithm fails to identify one line for thioformaldehyde (H2CS). The reason is that our method identified three alternative lines as illustrated in Figure 10.
@@ 4.2.3 OrionKL-Tianma
In terms of OrionKL-Tianma, we demonstrate the best-fitting spectra and the recalls in Figure 11 and Figure 12 respectively. The best-fitting spectrum using $\chi^{2}$ shows considerable wrong features, and a potential reason is that the observed spectrum consists of over 100 strong radio recombination lines, which disturbs the fitting. In contrast, the result using the peak matching loss function well agrees with the observed spectrum, indicating the robustness of our method.
The result using the peak matching loss function only identifies about half acetone (CH3COCH3) lines, missing 9 lines. In fact, our method reproduces the right features, while these features are slightly below the threshold of our peak finding algorithm. Consequently, these features are not counted in computing the recall. We demonstrate this point in Figure 13. The similar problem is also found in the results of CH3OCHO, C2H5CN, and 34SO2.
@@
Figure 14 compares the parameters estimated by Liu et al. (2022) with our best-fitting parameters. Whenever the original study suggested multiple components, the component with the highest column density is compared. For each molecule, Liu et al. (2022) used fixed values for the source size and excitation temperature, and fit the column density, velocity width, and velocity offset. In Figure 14, for the result based on the peak matching loss, we find that the column density and velocity offset are generally consistent, whereas the source size, excitation temperature, and velocity width are inconsistent. The result using the $\chi^{2}$ loss tends to predict larger velocity widths.
@@ 4.2.4 SgrB2M-IRAM
We show the best-fitting spectra for SgrB2M-IRAM in Figure 15, and the recalls are given in Figure 16. While the result using the $\chi^{2}$ loss function achieves an overall recall of 60%, the best fitting spectrum is fundamentally flawed. In contrast, in such line rich region, our result using the peak matching loss fits the observed spectrum well, with an overall recall of 85 %.
The result using the peak matching shows a low recall of ethanol (C2H5OH), which is mainly due to line blending. To illustrate the point, we show the best fitting spectrum of C2H5OH in the individual fitting phase as blue lines, which can match most of the lines. However, by comparing the result obtained in the individual fitting phase with the final spectrum, we can find line blending at 82650.97 MHz, 102487.53 MHz, 106648.67 MHz, and 114066.018 MHz.
Our results also attain low recalls of HCCCN and CH3OH. These molecules tend to be better described by multiple physical components, while our method assumes one component for each molecule for simplicity. This issue is also mentioned in Section 4.2.1.
In Figure 18, we compare the estimated parameters from Belloche et al. (2013) with our results. Whenever the original study suggests multiple components, the component with the highest column density is compared. Belloche et al. (2013) also generated the spectrum using xclass. While a correlation is found for the column density, our results based on the peak matching loss tend to underestimate it by 1 - 2 dex. A potential reason could be that our algorithm tends to adopt larger source sizes than Belloche et al. (2013). Moreover, no significant correlation can be found for the other parameters.
@@ 4.2.5 SgrB2N-IRAM
[…]
The identification of cyclopropenylidene (c-C3H2) is special, which is affected by absorption in the observed spectrum. Figure 21 illustrates the spectrum of c-C3H2 generated using the parameters suggested by Belloche et al. (2013). Absorption features are found at the locations of three peaks, at 82093.63MHz, 85339.04MHz, and 87435.74 MHz. If the absorption component is modeled properly, these three peaks will be attenuated. However, this work does not model absorption, and the absorbed spectrum is simply set to zero. Consequently, the three peaks are treated as false positive peaks in our algorithm, resulting in a higher loss. Whereas our result identifies the spectral lines of c-C3H2, our algorithm cannot handle the case associated with absorption in general.
In addition, our algorithm fails to identify HNCS, CH3COOH, and HNC18O. We find it difficult to reproduce the spectral lines of these molecules using the parameters suggested by Belloche et al. (2013). A potential reason could be that different spectral data were used.
In Figure 22, we compare the parameters suggested by Belloche et al. (2013) with our best-fitting parameters. Both results employ xclass to generate model spectra. Similar to SrgB2M-IRAM, we can only find correlation for the column density, and the other parameters show discrepancies.
@@ 4.3 Computational cost
For all line surveys considered in this study, the automated line identification algorithm is executed using 28 2.6 GHz CPUs in parallel. When employing the peak matching loss function, the time spent on each line survey is given in Table 2. The computational cost varies from 5.6 to 59.0 hours, depending on the complexity of the observed spectrum.
@@ 5 Summary
  • We propose the peak matching loss function, which significantly enhances the robustness of spectral fitting. Our results using the peak matching loss are consistently better than using the widely used $\chi^{2}$ loss. It is worth noting that the peak matching loss function is an independent component in our method, and in principle can be incorporated into other codes to improve the fitting results.
@@
We evaluate our algorithm using published data of five line surveys, including line rich regions in Sgr B2. Our algorithm achieves an overall recall of 84% - 98%. The errors of our algorithm are mainly due to the following factors:
@@
We also compare our best-fitting parameters with the previous studies. Whereas our best-fitting spectra are in good agreement with the observations, the estimated parameters are in general inconsistent with the previous studies. We find a correlation only for the column density, with an error of two orders of magnitude. We attribute this discrepancy to the inherent degeneracy between the source size, excitation temperature, and column density.
@@
Figure 6: Recall of molecules identified in W51-Mopra. Red and blue bars illustrate the results using the peak matching and $\chi^{2}$ loss functions respectively. The recall, the number of matched lines, and the number of relevant lines are indicated on top of each bar. Refer to Section 4.1 for the definition of recall.
Figure 7: Spectral lines of methanol in W51-Mopra. Black solid lines show the observed spectrum from Watanabe et al. (2017), and black vertical lines are spectral lines of methanol identified by Watanabe et al. (2017). Red and blue lines show the fitting results using one and two physical components respectively. The recall of each fitting result is indicated in the first panel. These plots imply that the spectrum of methanol can be better modeled using two physical components.
@@
Figure 9: Recall of molecules identified in OrionKL-GBT. Red and blue bars illustrate the results using the peak matching and $\chi^{2}$ loss functions respectively. The recall, the number of matched lines, and the number of relevant lines are indicated on top of each bar. Refer to Section 4.1 for the definition of recall.
Figure 10: Identified lines of H2CS in OrionKL-GBT. The result is obtained using the peak matching loss function. The result implies a different component of H2CS in addition to the one identified by Frayer et al. (2015).
@@
Figure 12: Recall of molecules identified in OrionKL-Tianma. Red and blue bars illustrate the results using the peak matching and $\chi^{2}$ loss functions respectively. The recall, the number of matched lines, and the number of relevant lines are indicated on top of each bar. Refer to Section 4.1 for the definition of recall.
Figure 13: Line identification of acetone (CH3COCH3) in OrionKL-Tianma. Black solid lines show the observed spectrum. Black dashed lines indicate the CH3COCH3 spectral lines identified by Liu et al. (2022). Grey horizontal lines represent the threshold used by our algorithm to identify a peak. Red lines show the best-fitting spectrum using the peak matching loss. While the best-fitting spectrum predicts most peaks at the identified lines of Liu et al. (2022), some peaks are too faint to be identified by our peak finder. This explains the low recall of CH3COCH3, as shown in Figure 12.
Figure 14: Comparison of the parameters estimated by Liu et al. (2022), labeled ”test”, with the best-fitting parameters derived from the peak match loss function, labeled ”pred”. Whenever the original study suggests multiple components, the component with the highest column density is compared. Red and blue dots represent the results based on the peak matching and $\chi^{2}$ loss functions respectively. Each data point compares the parameters of a molecule at one state. Liu et al. (2022) adopted fixed values for source size and excitation temperature, and fit the column density, velocity width, and velocity offset.
@@
Figure 16: Recall of molecules identified in SgrB2M-IRAM. Red and blue bars illustrate the results using the peak matching and $\chi^{2}$ loss functions respectively. The recall, the number of matched lines, and the number of relevant lines are indicated on top of each bar. Refer to Section 4.1 for the definition of recall.
Figure 17: Line blending analysis of ethanol (C2H5OH) in SgrB2M-IRAM. Black solid lines represent the observed spectrum, and black dashed lines indicate the C2H5OH spectral lines identified by Belloche et al. (2013). Blue lines show the best-fitting spectrum of C2H5OH obtained in the individual fitting phase. Red solid lines illustrate the spectrum of C2H5OH as a part of the combined spectrum. The final combined spectrum is shown as red dashed lines. While the final result show a low recall of C2H5OH, most lines can be recovered in the individual fitting phase, implying line blending.
@@
Figure 20: Recall of molecules identified in SgrB2N-IRAM. Red and blue bars illustrate the results using the peak matching and $\chi^{2}$ loss functions respectively. The recall, the number of matched lines, and the number of relevant lines are indicated on top of each bar. Refer to Section 4.1 for the definition of recall.
Figure 21: Line identification of cyclopropenylidene (c-C3H2) by Belloche et al. (2013). Black solid lines represent the observed spectrum, and black dashed lines indicate the spectral lines of c-C3H2 identified by Belloche et al. (2013). Purple lines represent the spectrum generated using the parameters suggested by Belloche et al. (2013). Each panel displays a peak from the model spectrum.
This work was supported by the National Natural Science Foundation of China under grant 12373026, and the Leading Innovation and Entrepreneurship Team of Zhejiang Province of China (Grant No. 2023R01008). The computation is performed at the Shuguang supercomputer in Zhejiang Laboratory.

### v2

@@ 4 Applying to line surveys
In this study, for speed reason, we downsample all observed spectra by averaging every two adjacent channels. The downsampling is repeated three times. Accordingly, the number of channels in the observed spectra is reduced by a factor of eight, and the RMS noise values are divided by $\sqrt{8}$ before being input into the peak finding algorithm. Secondly, for simplicity, this work neglects any absorption in the observed spectra. Given that the baselines of all the observed spectra used in this work have been calibrated to zero by the original studies, our correction is straightforward. Specifically, any channels displaying a negative value are reset to zero. In addition, the continuum spectra are not taken into account in the fitting processes. This approach may have negligible effects if the excitation temperature of the molecule is significantly higher than that of the continuum. Conversely, for molecules whose excitation temperature is comparable to the background temperature, neglecting the continuum spectra could affect the estimated excitation temperature and column density.
Furthermore, our algorithm requires a molecule list as input. Since this work focuses on whether our method can reproduce the identification results of the previous studies rather than identifying new molecules, we use the union of the molecules and isotopologues identified in all the five line surveys as the input list, which contains 105 species. For practical use, our code provides a list of commonly observed molecules, e.g. methanol (CH3OH), methyl formate (CH3OCHO) and cyanoacetylene (HCCCN). Users can customize the molecule list to suit their specific applications. Users may also refer to McGuire (2022) for molecules that have been detected in different sources.
@@ 4.1 Metrics
In this work, we use recall and precision to compare our identification results with the published line survey data. Recall measures the ability of our algorithm to find all spectral lines identified by the previous studies, while precision measures the accuracy of our line assignments.
To compute the metrics, for each line survey, we prepare a line table for comparison that contains only lines identified by our peak finder in both the final predicted spectrum and the downsampled observed spectrum. The downsampling introduced earlier may result in deviations of the peak frequencies. To match the lines between the predicted and observed spectra, we use the peak widths determined by the scipy function peak_widths (see Section 2.4) as the tolerance criteria. When constructing the line tables for comparison, the following points are taken into account:
@@
    A peak in the observed spectrum may be labeled by multiple transitions from the same molecule. In such cases, these transitions are treated as a single line.
@@
    If a line in the observed spectrum is labeled by multiple different species, that line is excluded. Since the original study does not provide the contributions for each species, comparing such lines is tricky. In contrast, if a line in the predicted spectrum is labeled by multiple different species, we label the line with the one that has the largest contribution.
@@
    The line surveys may use different spectral line databases, and to be consistent, we only include lines that are present in the spectral line database used in this work.
@@
    Radio recombination lines identified in the observed spectrum are excluded.
We define the recall and precision metrics as
Recall $=\frac{N_{\text{match}}}{N_{\text{ref}}},$ (27)
Precision $=\frac{N_{\text{match}}}{N_{\text{pred}}}.$ (28)
For a given molecule, $N_{\text{ref}}$ is the number of lines identified by the referred study, $N_{\text{pred}}$ is the number of lines labeled as that molecule by our algorithm, and $N_{\text{match}}$ is the number of lines that match between the referred study and our algorithm. For example, if the referred study identifies 10 methanol lines, and our algorithm labels 9 methanol lines, where 8 lines match those in the referred study, and 1 line is labeled as a different molecule, then the recall of methanol is $8/10=0.8$, and the precision of methanol is $8/9\approx 0.89$. In addition, when counting $N_{\text{ref}}$, $N_{\text{pred}}$, and $N_{\text{match}}$, we only consider those in the line table for comparison as described in the previous paragraph.
Furthermore, we compute the overall recall for each line survey by dividing the total number of lines identified by our algorithm by the number of lines in the table for comparison. We also calculate the average precision for each line survey, which is the average of the precisions for all molecules.
@@ 4.2 Results
Table 3 presents a comparison of the overall recall and average precision obtained using the $\chi^{2}$ and peak matching loss functions for each line survey. Across all the line survey data, the peak matching loss function consistently outperforms the $\chi^{2}$ loss function in terms of both the recall and precision values.
We illustrate the issue with using the $\chi^{2}$ loss function through an example. In Figure 4, we present the fitting results of C2H3CN during the individual fitting phase for SrgB2N-IRAM. The results based on the peak matching and $\chi^{2}$ loss functions are represented by red and blue lines respectively. It is evident that the red line gives a better result than the blue line. However, upon calculating the $\chi^{2}$ and peak matching loss values for both fits, we find that the $\chi^{2}$ value of the blue line ($\chi^{2}$ = 12780.8) is smaller than the red line ($\chi^{2}$ = 12781.0). In other words, if the $\chi^{2}$ loss is adopted, the algorithm will consider the blue line as a better result than the red line, irrespective of optimization methods. Such result is unexpected. In contrast, the proposed peak matching loss assigns a lower value to the red line ($\mathcal{L}_{\text{PM}}=-2.81$) than the blue line ($\mathcal{L}_{\text{PM}}=0.89$), which appropriately reflects the variance between the observed and fitting spectra.
The results discussed above align with the practice of using $\chi^{2}$, where careful settings of initial guess and search space are necessary. For example, the result based on the $\chi^{2}$ loss can be improved by imposing an upper limit of the line width. These findings imply that the $\chi^{2}$ loss function may not be ideal for an automated spectral line fitting framework, where we aim for minimal manual settings, whereas the proposed peak matching loss function appears to be a superior alternative.
@@ 4.2.1 W51-Mopra
We illustrate the best-fitting spectrum for W51-Mopra in Figure 5, and show the recalls and precisions in Figure 6. The overall recall improves from 61% using the $\chi^{2}$ loss to 84% using the peak matching loss. The average precision also increases from 82% to 93%. The recall of methanol stands relatively low, at 54%. Our analysis suggests that the spectrum of methanol can be better described by more than one component. As demonstrated in Figure 7, by adding an additional component, the best-fitting spectrum of methanol is improved, and the recall increases to 96%.
@@ 4.2.2 OrionKL-GBT
For OrionKL-GBT, we present the best-fitting spectra in Figure 8. The recall and precision metrics of the molecules are demonstrated in Figure 9.Frayer et al. (2015) identified only those lines that are brighter than 1 K, which is significantly above the RMS noise. Therefore, we obtain the best results on OrionKL-GBT, with a recall of 93 % using the peak matching loss. Our algorithm fails to identify one line for thioformaldehyde (H2CS). Our analysis suggests that the emission from thioformaldehyde involves two velocity components. Besides the one line identified by Frayer et al. (2015), our method has detected a different velocity component comprising three lines, as illustrated in Figure 10. Additionally, our algorithm misidentified isocyanic acid (HNCO), resulting from an incorrect assignment. As shown in Figure 11, the two HNCO lines identified by Frayer et al. (2015) were incorrectly assigned to HN13CO in our result.
@@ 4.2.3 OrionKL-Tianma
In terms of OrionKL-Tianma, we demonstrate the best-fitting spectra in Figure 12, the recall and precision metrics in Figure 13. The best-fitting spectrum using $\chi^{2}$ shows considerable wrong features, and a potential reason is that the observed spectrum consists of over 100 strong radio recombination lines, which disturbs the fitting. In contrast, the result using the peak matching loss function well agrees with the observed spectrum, indicating the robustness of our method.
The result using the peak matching loss function only identifies less than half acetone (CH3COCH3) lines, missing 11 lines. We demonstrate the issue in Figure 14. In fact, our method reproduces the right features, while these features are slightly below the threshold of our peak finding algorithm. Consequently, these features are not counted in computing the recall. The similar problem is also found in the results of CH3OCHO, C2H5CN, and 34SO2.
@@
The upper panel of Figure 15 compares the parameters estimated by Liu et al. (2022) with our best-fitting parameters. When the original study suggested multiple components of a molecule, we compare the component with the highest integrated intensity. In Figure 15, for the result based on the peak matching loss, we find that the column density and velocity offset are generally consistent, whereas the source size, excitation temperature, and velocity width are inconsistent. For each molecule, Liu et al. (2022) fit only the column density, velocity width, and velocity offset, while used fixed values for the source size and excitation temperature, which could account for the discrepancies in the source size and excitation temperature.
In the lower panel of Figure 15, we compare the fitting parameters obtained with the peak matching loss and the $\chi^{2}$ loss functions. The predicted values for the column density, and velocity offset are correlated, while the other parameters display significant variability. In particular, the results obtained with the $\chi^{2}$ loss tends to predict broader velocity widths.
@@ 4.2.4 SgrB2M-IRAM
For SgrB2M-IRAM, Figure 16 shows the best-fitting spectra, and Figure 17 illustrates the recall and precision values for each molecule. For the results based on the $\chi^{2}$ loss function, the best-fitting spectrum is fundamentally flawed, with a overall recall of 39% and an average precision of 56%. In contrast, for this line rich region, our result using the peak matching loss fits the observed spectrum well, achieving an overall recall of 78% and an average precision of 92%.
Our results show a low recall of ethanol (C2H5OH) at 38%, potentially due to line blending. To demonstrate the problem, we present the best-fitting results using the peak matching loss in Figure 18. The blue lines represent the best-fitting spectrum of C2H5OH from the individual fitting phase, aligning with most of the C2H5OH lines (indicated by black vertical lines) as identified by Belloche et al. (2013). However, these lines are absent in the final combined spectrum, implying that the fitting spectrum of C2H5OH is influenced by contributions from other species.
In Figure 19, we compare the estimated parameters from Belloche et al. (2013) with our results. When the original study suggested multiple components of a molecule, we compare the component with the highest integrated intensity.Belloche et al. (2013) also generated the spectra using xclass, and their parameters were adjusted through visual comparison with the observed spectrum. While a correlation is found for the column density, our results based on the peak matching loss tend to underestimate it by 1 - 2 dex. A potential reason could be that our algorithm tends to adopt larger source sizes than Belloche et al. (2013). Moreover, no significant correlations can be found for the other parameters. In particular, the predicted $\delta v_{\text{LSR}}$ values exhibit large uncertainties. We estimated that the standard deviation of the velocity offset errors, $v^{\text{test}}_{\text{LSR}}-v^{\text{pred}}_{\text{LSR}}$, is $\sim$ 3.0 km/s, larger than the standard deviation of $v^{\text{test}}_{\text{LSR}}$, at 1.0 km/s. This could explain the absence of a correlation between $v^{\text{test}}_{\text{LSR}}$ and $v^{\text{pred}}_{\text{LSR}}$. However, we consider the errors in the predicted velocity offsets to be acceptable, since the mean velocity width of the test samples is 12.4 km/s, which is four times the standard deviation of the velocity offset errors. Additionally, we verified that the significant deviations of the predicted $\delta v_{\text{LSR}}$ result from poor fitting. For instance, as previously discussed above, the metrics for C2H5OH are unsatisfactory. The predicted $\delta v_{\text{LSR}}$ for C2H5OH is 5.8 km/s, whereas Belloche et al. (2013) reported a value of 0.0 km/s.
@@ 4.2.5 SgrB2N-IRAM
SrgB2N-IRAM is the most challenging test data in this study, characterized by a line density of 59.2 GHz-1. In Figure 20, the best-fitting spectrum based on the peak matching loss is in good agreement with the observed spectrum, while the outcome derived from the $\chi^{2}$ loss is inconclusive. Regarding the metrics depicted in Figure 21, the peak matching loss function demonstrates superior performance compared to the $\chi^{2}$ loss function, with the overall recall rising from 31% to 74%, and the average precision from 51% to 78%.
The identification of cyclopropenylidene (c-C3H2) is a special case, affected by absorption in the observed spectrum. Figure 22 displays the c-C3H2 spectrum generated using the parameters suggested by Belloche et al. (2013). Absorption features are found at the locations of three peaks, at 82094 MHz, 85339 MHz, and 87436 MHz. Proper modeling of the absorption component would lead to the attenuation of these three peaks. However, in this study, absorption is not taken into account, and the affected spectrum is set to zero. As a result, these three peaks are classified as false positives by our algorithm, leading to an increased loss. Consequently, the identification c-C3H2 becomes challenging for our algorithm.
In addition, our algorithm fails to identify HNCS and HNC18O. We find it difficult to reproduce the spectral lines of these molecules using the parameters suggested by Belloche et al. (2013). A potential reason could be that different spectral data were used.
In Figure 23, we compare the parameters suggested by Belloche et al. (2013) with our best-fitting parameters. Both results employ xclass to generate model spectra. Similar to SrgB2M-IRAM, we can only find correlation for the column density, and the other parameters show discrepancies. The potential reasons have been discussed in Section 4.2.4.
@@ 4.3 Computational cost
Figure 24 demonstrates the computation time of our automated line identification method with the peak matching loss function. For all line surveys analyzed in this study, our algorithm is executed using 28 2.6 GHz CPUs in parallel. The computational cost varies from 26 to 91 hours, depending on the complexity of the observed spectrum. The computation resources required by our method are relatively demanding. We find that the majority of the computational time is devoted to calling xclass, while the time allocated to optimization algorithm itself and the calculation of the peak matching loss function is minimal. Accordingly, enhancing the efficiency of the spectral line model could expedite our line identification pipeline.
@@ 5 Summary
  • We have shown that the widely used $\chi^{2}$ loss function tends to produce unsatisfactory results under minimal human supervision. To address this issue, we propose the peak matching loss function, which significantly enhances the robustness of spectral fitting. When applying to our automated line identification framework, the results obtained with the peak matching loss function are consistently better than those from the $\chi^{2}$ loss function. It is worth noting that the peak matching loss function is an independent component in our method, and in principle can be incorporated into other codes to improve the fitting results.
@@
We evaluate our algorithm using published data of five line surveys, including line rich regions in Sgr B2. Our algorithm achieves an overall recall of 74% - 93% and an average precision of 78% - 92%. The errors of our algorithm are mainly due to the following factors:
@@
We also compare our best-fitting parameters with the previous studies. Whereas our best-fitting spectra are in good agreement with the observations, the estimated parameters are in general inconsistent with the previous studies. We find a correlation only for the column density, with an error of two orders of magnitude. We attribute this discrepancy to the inherent degeneracy between the source size, excitation temperature, and column density. We suggest that the uncertainties of the fitting parameters can be better studied using mock data where the true parameters are known. Such studies could constitute future work.
@@
Figure 6: Recall and precision of species identified in W51-Mopra. The red and blue bars represent the results using the peak matching and $\chi^{2}$ loss functions respectively. Definitions of recall and precision are provided in Section 4.1.
Figure 7: Spectral lines of methanol in W51-Mopra. The black solid lines show the observed spectrum from Watanabe et al. (2017), and the black vertical lines are spectral lines of methanol identified by Watanabe et al. (2017). The red and blue lines show the fitting results using one and two physical components respectively. The recall of each fitting result is indicated in the legend. These plots imply that the spectrum of methanol can be better modeled using two physical components.
@@
Figure 9: Recall and precision of species identified in OrionKL-GBT. The red and blue bars represent the results using the peak matching and $\chi^{2}$ loss functions respectively. Definitions of recall and precision are provided in Section 4.1.
Figure 10: Identified lines of H2CS in OrionKL-GBT. The red lines show our best-fitting spectrum of H2CS using the peak matching loss function. The result implies a different velocity component of H2CS in addition to the one identified by Frayer et al. (2015).
Figure 11: Incorrect line assignment of HNCO in OrionKL-GBT by our algorithm. The black vertical dashed lines indicate the two HNCO lines identified by Frayer et al. (2015). Our algorithm incorrectly assign them to HN13CO as indicated by the red lines, which show the best-fitting spectrum of HN13CO using the peak matching loss function.
@@
Figure 13: Recall and precision of species identified in OrionKL-Tianma. The red and blue bars represent the results using the peak matching and $\chi^{2}$ loss functions respectively. Definitions of recall and precision are provided in Section 4.1.
Figure 14: Line identification of acetone (CH3COCH3) in OrionKL-Tianma. The black solid lines show the observed spectrum, and the black dashed lines indicate the CH3COCH3 spectral lines identified by Liu et al. (2022). The gray horizontal lines represent the threshold used by our algorithm to identify a peak. The red lines show the best-fitting spectrum using the peak matching loss. While the best-fitting spectrum predicts most peaks at the identified lines by Liu et al. (2022), some peaks are too faint to be identified by our peak finder. This explains the low recall of CH3COCH3, as shown in Figure 13.
Figure 15: Upper panels: comparison of the parameters estimated by Liu et al. (2022), labeled ”test”, with our best-fitting parameters, labeled ”pred”. When the original study suggests multiple components, we compare the component with the highest integrated intensity. The red and blue dots represent the results based on the peak matching and $\chi^{2}$ loss functions respectively. Each data point compares the parameters of a molecule at one state. Liu et al. (2022) adopted fixed values for source size and excitation temperature, and fit the column density, velocity width, and velocity offset. Lower panels: comparison of the parameters derived using the peak matching and $\chi^{2}$ loss functions.
@@
Figure 17: Recall and precision of species identified in SgrB2M-IRAM. The red and blue bars represent the results using the peak matching and $\chi^{2}$ loss functions respectively. Definitions of recall and precision are provided in Section 4.1.
Figure 18: Line blending analysis of ethanol (C2H5OH) in SgrB2M-IRAM. The black solid lines represent the observed spectrum, and the black dashed lines indicate the C2H5OH spectral lines identified by Belloche et al. (2013). The blue lines show the best-fitting spectrum of C2H5OH obtained in the individual fitting phase. The red solid lines illustrate the spectrum of C2H5OH as a part of the combined spectrum. The final combined spectrum is shown as the yellow lines. While the final result show a low recall of C2H5OH, most lines can be recovered in the individual fitting phase, implying line blending.
Figure 19: Comparison of the parameters estimated by Belloche et al. (2013), labeled ”test”, with the best-fitting parameters obtained in this work, labeled ”pred”. Whenever the original study suggests multiple components, the component with the highest column density is compared. Each data point compares the parameters of a molecule at one state. The red and blue dots represent the results based on the peak matching and $\chi^{2}$ loss functions respectively.
Figure 21: Recall and precision of species identified in SgrB2N-IRAM. The red and blue bars represent the results using the peak matching and $\chi^{2}$ loss functions respectively. Definitions of recall and precision are provided in Section 4.1.
Figure 22: Line identification of cyclopropenylidene (c-C3H2) by Belloche et al. (2013). The black solid lines represent the observed spectrum, and the black dashed lines indicate the spectral lines of c-C3H2 identified by Belloche et al. (2013). The purple lines represent the spectrum generated using the parameters suggested by Belloche et al. (2013). Each panel displays a peak from the model spectrum. These plots imply that absorption should be taken into account when fitting the c-C3H2 spectrum.
@@
Figure 24: Computation time for our automated line identification algorithm using the peak matching loss function.
acknowledgments
This work was supported by the National Natural Science Foundation of China (Grant No. 12373026), the Leading Innovation and Entrepreneurship Team of Zhejiang Province of China (Grant No. 2023R01008), the Key R&D Program of Zhejiang, China (Grant No. 2024SSYS0012), and Young Scientists Fund of the National Natural Science Foundation of China (Grant No. 12403030). The computation is performed at the Shuguang supercomputer in Zhejiang Lab. We thank the reviewer for providing a constructive report, which improved the quality of the paper.